OpenStudy (anonymous):

A ball is rolled over a level lawn with initial velocity 8ms^-1. due to the friction the velocity decreases and is given by dv/dt=-2. Find the equation of distance travelled?

OpenStudy (anonymous):

First integrate dv/dt =-2 so it becomes v = -2t +C, at t =0, v=8, plug in, find C then that's the equation of velocity. then integrate velocity equation

OpenStudy (anonymous):

so, what the meaning of equation of distance travelled? add the equation we get or what?

OpenStudy (anonymous):

it's the equation of distance travel in term of t, like s(t)= ...., and when you integrate velocity equation, you get distance equation

OpenStudy (anonymous):

So, s(t) = -t^2 +8t +C

OpenStudy (anonymous):

And I think at time t=0, the distance travel is 0, so 0= 0+0 +C, so C=0 s(t) = -t^2 +8t is distance equation

OpenStudy (anonymous):

does this answer your question ?

OpenStudy (anonymous):

why we must integrate the velocity equation again?

OpenStudy (anonymous):

At first, you're given dv/dt, which is the derivative of velocity, or acceleration. After you integrate it, it becomes velocity equation, which is v(t) = -2t +8. THen you integrate again to get position equation or distance travel.

OpenStudy (anonymous):

oo i understand now..k tq.. :)